A human sperm is about 14x10^3 mm long and can travel 15cm in 3 hours.

1) What distance does the sperm swim in relation to its body length. (2 marks)

2) Calculate the distance a 1,8m male would swim if he could swim as far as his sperm, compared with his own height (4 marks)
1
Expert's answer
2017-06-02T12:16:09-0400
1. Distance the sperm swim in relation to its body length: D 1 = 150 mm / 14x10 -3 = 10714.
2. Distance a 1,8 m male would swim if he could swim as far as his sperm, compared with his own height: D 2 = 1.8 · 10714 = 19258 m.